### v1

# CodeAssistBench (CAB): Dataset & Benchmarking for Multi-turn Chat-Based Code Assistance

Abstract
Programming assistants powered by large language models have transformed software development, yet most benchmarks focus narrowly on code generation tasks. Recent efforts like InfiBench and StackEval attempt to address this gap using Stack Overflow data but remain limited to single-turn interactions in isolated contexts, require significant manual curation, and fail to represent complete project environments. We introduce CodeAssistBench (CAB), the first benchmark framework for evaluating multi-turn programming assistance in realistic settings that address real-world questions about actual codebases. Unlike existing programming Q&A benchmarks, CAB automatically generates scalable datasets from question-related GitHub issues using configurable parameters (e.g., repository creation date, star count, programming languages), and includes automatic containerization of codebases for evaluation. It then evaluates models through simulated users in these containerized environments with full codebase access. Using this framework, we constructed a test set of 3,286 real-world programming questions across 231 repositories, spanning seven programming languages and diverse problem domains. Our evaluation of leading LLMs reveals a substantial capability gap: while models perform well on Stack Overflow questions with success rates of 70-83%, they resolve only up to 16.49% of CAB’s recent issues. This discrepancy highlights the challenges of providing assistance in complex, project-specific contexts versus answering standalone questions. Our fully automated framework enables continuous benchmark expansion and is available at https://anonymous.4open.science/r/CAB-CBA3/.
1 Introduction
Large language models (LLMs) are increasingly used throughout programming workflows. Early benchmarks such as HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) focused on generating functionally correct code snippets for isolated problems. As the field matured, benchmarks like CodeContests Li et al. (2022) introduced harder tasks; while others, including SWE-benchh Jimenez et al. (2023), BigCodeBench Zhuo et al. (2024), and SWE-PolyBench Rashid et al. (2025) began to address more realistic scenarios. Multi-turn benchmarks such as ConvCodeWorld Han et al. (2025), MINT Wang et al. (2023), and TICODER Fakhoury et al. (2024) further advanced this direction by incorporating conversational interactions, though they still primarily evaluate code synthesis. Despite these advances, most benchmarks fail to capture the full spectrum of developer needs. A recent Stack Overflow survey of 34,168 developers reveals that when asked about future AI tool expectations, developers expressed strong interest across multiple assistance categories. While 75.7% want AI for code generation, even higher percentages desire AI support for searching for answers (77.9%), debugging and troubleshooting (77.3%), and learning about codebases (73.6%) Stack Overflow (2024). These findings demonstrate that developers anticipate using AI tools as comprehensive collaborators throughout the entire development lifecycle.
Recent benchmarks like InfiBench Li et al. (2024) and StackEval Shah et al. (2024) incorporate Stack Overflow data to better reflect developer Q&A scenarios. However, these benchmarks have critical limitations: they use only single-turn interactions in isolated contexts, require extensive manual curation, and fail to represent real project environments where developers engage in iterative exchanges - trying suggestions, encountering errors, and requesting clarifications. The text clearly states that current language models are achieving high success rates on two benchmarks (InfiBench at 70.64% and StackEval at 83.0%), and shows that these benchmarks no longer represent significant challenges for recent LLMs.
Figure 1: Port conflict clarification in a real-world exchange.
Most multi-turn benchmarks for code generation have been limited to programming tasks, where evaluation typically relies on a binary oracle - running tests to check if they pass. Figure 1, parsed from a real GitHub issue, shows why that criterion fails for programming assistance. Even a trivial port-mapping question forces an assistant to (1) infer project topology, (2) explain that the proxy port is hard-coded, and (3) reassure the user that no extra mapping is needed. Each reply reshapes the next user utterance, so we must judge an evolving conversation, not a single execution. What matters is a gradient of clarity, efficiency, and user confidence that unfolds over turns. Evaluating it demands benchmarks that track satisfaction conditions across the dialogue, not just test outcomes.
To address these limitations, we introduce CodeAssistBench (CAB), a chat-based benchmark framework for evaluating multi-turn programming assistance grounded in real-world developer scenarios. Our contributions include:
1. A fully automated data generation framework that enables continuous benchmark creation and expansion without manual curation. This automation allows us to create dataset variants including recent repositories (post-training cutoff), ensuring the benchmark remains challenging as models evolve.
2. A realistic multi-turn Q&A dataset with 3,286 real-world programming questions across 231 repositories, spanning seven programming languages and diverse problem domains, all derived from actual GitHub issues.
3. A novel chat-driven evaluation pipeline that simulates realistic user-agent interactions in containerized environments with full codebase access. Our pipeline includes a simulated user that provides contextual feedback, a maintainer agent with environment access, and an automated judge that evaluates conversation quality against extracted satisfaction conditions.
Our evaluation of leading LLMs reveals a substantial capability gap: even the best model solve only 16.49% of recent CAB issues compared to 29.14% on historical data. This performance difference highlights both the challenges of CAB and the value of our framework in generating continuously evolving benchmarks. To advance AI coding assistance, we release the complete CAB benchmark, evaluation framework, and detailed results publicly here.
2 Related Work
Although programming assistance spans diverse domains, current benchmarks fail to reflect the full complexity of real-world developer support.
Code Generation Benchmarks. Traditional evaluations of programming assistants have focused on functional code generation. HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and CodeContests Li et al. (2022) assess model ability to generate functionally correct code from well-specified prompts. More advanced benchmarks such as SWE-Bench Jimenez et al. (2023) and SWE-PolyBench Rashid et al. (2025) use real GitHub issues to evaluate software maintenance tasks across actual repositories. While these provide strong correctness metrics using pass@1, they are largely limited to single-shot interactions and do not capture the iterative nature of real developer workflows.
Multi-turn Programming and Conversational Benchmarks. Several benchmarks aim to evaluate iterative interactions. ConvCodeWorld Han et al. (2025), MINT Wang et al. (2023), and TICODER Fakhoury et al. (2024) simulate developer-assistant dialogues but typically assume consistent environments and remain centered on code synthesis. More general conversational benchmarks—such as MT-Bench Zheng et al. (2023), AlpacaEval Dubois et al. (2023), and BotChat Duan et al. (2024) - offer valuable insights into dialogue quality but do not target the unique demands of technical problem-solving in programming contexts. In contrast, our work directly addresses this intersection by evaluating LLMs in multi-turn, task-grounded developer scenarios across full codebases.
Programming Q&A Benchmarks. Recent benchmarks like StackEval Shah et al. (2024) and InfiBench Li et al. (2024) leverage StackOverflow data to better reflect how developers seek help in practice. StackEval evaluates LLMs on 925 questions across 25 languages and task types, while InfiBench offers 234 examples with model-agnostic evaluation metrics. These benchmarks better capture explanatory reasoning than traditional code benchmarks, but they are still constrained to single-turn formats and require significant manual curation, limiting scalability.
3 CAB: CodeAssistBench
Figure 2: Overview of CAB’s pipeline for dataset generation, consisting of: (1) Repository Collection from GitHub, filtering by stars, creation date, and license; (2) Filtering of issue conversations using regex and LLM-based criteria; and (3) Structured Data Generation including Build Environment, Satisfaction Conditions, and User Response References. This automated pipeline enables continuous benchmark expansion without manual curation.
CodeAssistBench (CAB) is a comprehensive benchmark framework designed to evaluate large language models (LLMs) in realistic, multi-turn programming assistance scenarios. Unlike prior benchmarks that focus on code generation or single-turn evaluation, CAB simulates developer workflows grounded in GitHub issues. Our framework consists of two key components: (1) an automated dataset generation pipeline that creates diverse, high-quality programming assistance scenarios from GitHub issues; and (2) a multi-agent evaluation framework that simulates realistic user-model interactions in containerized environments.
3.1 Dataset Generation Pipeline
The CAB dataset generation pipeline automatically transforms GitHub issues into structured programming assistance scenarios through three main stages: repository collection, issue filtering, and data preparation (Figure 2). This section elaborates on the details of this pipeline, while detailed statistics about the output of each phase are presented in Section 4.
3.1.1 Repository Collection
We begin by collecting a diverse set of GitHub repositories that represent real-world programming challenges across multiple languages. Let ${R}_{GH}$ denote the set of all public GitHub repositories. We extract a filtered subset:
$R=\left\{r\in\mathcal{R}_{GH}\mid s(r)>S_{\min},\;t(r)>t_{0},\;\ell(r)\in\mathcal{L}\right\}$ (1)
where:
  • $s(r)$ is the star count of repository $r$, with threshold $S_{\min}$ (e.g., 10),
  • $t(r)$ is its creation date, with cutoff $t_{0}$ (e.g., 2024-11-01),
  • $\ell(r)$ is $r$’s SPDX license, restricted to permissive licenses compatible with research use.
To prioritize repositories with active developer support communities, we define a community score:
$\mathrm{CS}(r)=Q(r)+H(r)$ (2)
where $Q(r)$ and $H(r)$ represents the number of closed issues labeled with “question” and “help wanted” respectively. We then select the top-$K$ repositories ranked by $\mathrm{CS}(r)$ (breaking ties by star count) to form our repository set $R_{K}$, where $K$ is a user-defined parameter.
3.1.2 Issue Filtering
For each repository $r\in R_{K}$, let $I_{r}$ be the set of issues associated with it. We fetch only successfully closed issues using the GitHub Issues API. Each issue is represented as:
$i=(\mathrm{title}(i),\,\mathrm{body}(i),\,\mathrm{messages}(i)),\quad\mathrm{messages}(i)=\{m_{i,1},m_{i,2},\dots\}$ (3)
We apply two filtering rules - implemented using regular expressions - to retain high-quality, interactive conversations: (1) requiring at least two distinct participants to ensure genuine question-answering dynamics; and (2) removing issues containing media content (e.g., URLs, images, videos) to focus on text-based programming assistance.
To ensure issue relevance and message quality, we apply two LLM-based filtering steps using structured prompts. At the issue level, we assess resolution status, specificity, clarity, and safety via a seven question prompt (see Listing A.1), asking the model to answer each of them with a binary Yes/No. Issues are retained only if they satisfy criteria such as being clearly resolved, technically specific, free of sensitive content (e.g., personal information), and reproducible. At the message level, we construct the full conversation and prompt the model to identify comments that provide no support-related value (e.g., "+1", "Thanks", "Bump"). Comments are preserved unless explicitly flagged for removal using strict exclusion rules (see Listing A.2).
After filtering, we restructure each issue as a sequence of turns, where a turn represents a complete user-maintainer interaction:
$\mathrm{turn}_{i,k}=(m_{i,k}^{\mathrm{author}},m_{i,k}^{\mathrm{maintainer}})$ (4)
The resulting filtered issue set $I^{\prime}_{r}$ contains issues with this turn-based structure:
$i=(\mathrm{title}(i),\,\mathrm{body}(i),\,\{\mathrm{turn}_{i,1},\mathrm{turn}_{i,2},\ldots\})$ (5)
3.1.3 Data Preparation
For each filtered issue, we prepare three essential components that enable realistic evaluation: build environment generation, user satisfaction condition extraction, and user response reference generation.
[…]
Satisfaction Condition Extraction. We use Sonnet3.7 with a structured prompt (provided in Appendix A.4) to identify explicit criteria that indicate when an issue is successfully resolved. The model analyzes the full conversation, focusing on the original question and subsequent clarifications, to extract concrete conditions that would satisfy the user’s needs. These form the set $s_{i}$ = $\{s_{i,1},\dots,s_{i,K}\}$ of satisfaction conditions that serve as objective evaluation criteria.
User Response Reference Generation. To enable realistic simulation of user follow-up behavior, we construct a BM25 index over historical maintainer-user message pairs (excluding data from the current issue) and retrieve the top-$N$ most similar maintainer responses with their corresponding user replies. These form the reference set $u_{i}$ = $\{u_{i,1},\dots,u_{i,M}\}$ that guides the simulated user’s feedback.
Result Aggregation. The resulting dataset entry for each issue is a tuple:
[…]
For each issue $i\in\bigcup_{r}I_{r}^{\prime}$, where $I_{r}^{\prime}$ denotes the filtered issues for repository $r$, we create such entries to form our complete dataset.
3.2 Evaluation Framework
Figure 3: CAB evaluation pipeline. A simulated user chats with the Maintainer Agent, which can run code in an optional build sandbox; the interaction continues until the user is satisfied or reaches the maximum turn limit. Once the dialogue ends, an LLM-judge grades the exchange against satisfaction conditions extracted from the original GitHub issue. This pipeline enables realistic assessment of programming assistance in context-rich, project-specific scenarios.
CAB’s evaluation framework simulates realistic programming assistance interactions through a multi-agent system with three distinct roles: a User, a Maintainer, and a Judge (Figure 3).
User Agent. The user agent initiates the conversation with a programming question extracted from a GitHub issue and provides follow-up responses based on the model’s answers. This agent presents the initial question with relevant context, evaluates model responses against satisfaction conditions, provides realistic follow-up questions or clarifications, and signals when a solution successfully resolves the issue. The user agent does not directly interact with the environment but observes execution results to assess progress toward resolution.
Maintainer Agent. The maintainer agent represents the LLM being evaluated. Given the user’s question and contextual information, the model must analyze the problem within the provided codebase context, execute commands in the containerized environment when necessary, generate helpful responses, and adapt its approach based on user feedback. To balance exploration with efficiency, the maintainer can take up to 5 exploration steps per response, where each step may involve file operations or command execution.
Judge Agent. After the conversation concludes, an automated judge evaluates the interaction quality based on: (1) technical correctness - whether the proposed solution correctly addresses the underlying issue; (2) satisfaction completeness - whether all extracted satisfaction conditions are met; and (3) interaction quality - whether the conversation was appropriately concise and helpful. For issues with Docker environments, execution success is a hard requirement - a technically sound explanation is insufficient if the implementation fails in practice. The evaluation is triggered when either the user expresses satisfaction or the conversation reaches a maximum of 10 turns. This approach provides a comprehensive assessment of an LLM’s ability to provide effective programming assistance in realistic scenarios.
3.3 Implementation Details
CAB is implemented as a fully automated pipeline that can continuously generate new benchmark datasets as GitHub repositories evolve. Our implementation integrates with GitHub’s REST API to collect repositories and issues with configurable parameters, automatically generates and validates Docker environments to ensure reproducibility, and orchestrates the multi-agent conversation flow while recording detailed interaction logs. This implementation enables CAB to generate continuously evolving benchmarks that remain challenging as models improve, particularly through the creation of dataset variants that include repositories created after model training cutoff dates.
4 Experimental Results and Analysis
In this section, we report statistics from our dataset generation pipeline, model evaluation experiments, and human alignment analysis. For details on the hardware setup and compute time associated with these experiments, we refer readers to Appendix D.
4.1 Statistics of Dataset Generation Pipeline
Table 1: Summary of issue filtering, Docker requirement detection, build outcomes, and final issue retention across programming languages and repository cohorts in our dataset generation pipeline.
Metric Total Python Java C++ C# JS TS C
Filtered Issues 3,342 660 543 518 545 445 443 188
   All-Time 3,033 500 535 487 520 406 417 168
   Recent 309 160 8 31 25 39 26 20
Docker-Required 294 90 58 53 30 34 12 17
   All-Time 252 64 58 48 28 30 9 15
   Recent 42 26 0 5 2 4 3 2
Successful Docker Builds 238 77 52 28 21 33 10 17
   All-Time 197 52 52 23 19 29 7 15
   Recent 41 25 0 5 2 4 3 2
Final Retained Issues 3,286 647 537 493 536 444 441 188
   All-Time 2,978 488 529 462 511 405 415 168
   Recent 308 159 8 31 25 39 26 20
We applied our dataset generation pipeline (Section 3.1) to two distinct repository cohorts - All-Time and Recent - to evaluate its scalability and effectiveness. The All-Time cohort comprises 700 top-starred repositories (100 per language) without creation date constraints, from which we selected 70 (10 per language) based on permissive licenses and community engagement. The Recent cohort includes 3,500 top starred repositories (500 per language) created after November 2024, from which 700 were selected using the same criteria - yielding a total pool of 770 repositories across seven programming languages  [^1] There is no overlap between the All-Time and Recent cohorts.
From these repositories, CAB collected 25,656 raw GitHub issues with question and help-wanted issue tags. A combination of regex - and LLM-based filters was used to discard issues with insufficient content or ambiguous resolutions. A summary of the filtered issues, along with Docker requirement classification, build results, and final issue retention, is presented in Table 1. Recent repositories exhibited a notably higher Docker build success rate - 97.6% (41 out of 42) - compared to 78.2% (197 out of 252) in All-Time repositories. Most build failures in older repositories stemmed from outdated dependencies or unavailable components, not from limitations of the LLM pipeline itself. To address this, CAB inferred project-specific requirements - such as the relevant library versions, operating systems, and toolchains - by analyzing both the issue’s content and its creation timestamp, and then tailored environment reconstruction accordingly via custom Dockerfile generation.
Ultimately, our dataset includes 3,286 multi-turn conversations from 214 repositories. Among these, 294 issues required Docker-based environments, and 238 were successfully built through our automated pipeline. A total of 44,628 Sonnet 3.7 calls were made across all steps, without any manual intervention after the initial setup. Language-wise, recent repositories skew toward Python and TypeScript, while All-Time repositories present a more balanced distribution.
As shown in Figure 5, over half of the issues across all languages involved multiple conversational turns. In particular, Python, C++, and TypeScript showed the highest rates of extended discussions (3+ turns), underscoring the need for multi-turn understanding and evaluation in real-world software development scenarios.
4.2 Model Evaluation Results
Table 2: Correctness and Average Turns for each model, based on both original GitHub threads and CAB-generated conversations. Rec refers to recent repositories, while All includes all-time repositories.
Model Correctness (%) Avg. GitHub Turn (Human) Avg. CAB Turn
Correct Partially Corr. Incorrect Correct Partially Corr. Incorrect Correct Partially Corr. Incorrect
Rec. All Rec. All Rec. All Rec. All Rec. All Rec. All Rec. All Rec. All Rec. All
ChatGPT 4.1 Mini 16.49 29.14 30.41 36.86 53.09 34.00 1.72 2.77 2.17 3.56 2.18 3.05 2.94 2.35 3.66 3.33 5.70 4.28
DeepSeek R1 11.34 27.14 33.51 40.29 55.15 32.57 2.05 2.66 2.32 3.62 1.98 2.99 2.82 2.24 3.18 2.72 4.50 4.28
Llama 3.3 70B 9.33 13.58 25.91 40.75 64.77 45.66 1.67 2.72 2.02 3.48 2.21 3.02 3.22 2.68 4.06 3.49 4.67 4.50
Haiku 3.5 7.22 16.86 30.93 43.43 61.86 39.71 1.79 2.63 1.93 3.47 2.23 3.04 3.86 2.73 3.97 3.81 6.76 5.63
Sonnet 3.7 11.34 25.71 30.93 35.43 57.73 38.86 1.68 2.86 1.95 3.45 2.27 3.09 2.36 2.30 3.57 3.15 5.71 4.21
Sonnet 3.7 (Think) 13.40 27.43 27.32 38.86 59.28 33.71 1.81 2.50 2.17 3.76 2.14 3.00 2.50 2.20 3.25 2.85 4.95 4.26
We benchmarked six state-of-the-art large language models (LLMs) - ChatGPT 4.1 Mini, DeepSeek R1, Llama 3.3 70B, Haiku 3.5, Sonnet 3.7, and Sonnet 3.7 (Think mode) - as maintainer agents using the CAB dataset. Each model engaged in multi-turn conversations to resolve user issues, guided by satisfaction conditions and evaluated by an automated assessor.
4.2.1 Metrics
We evaluate model performance using two primary metrics: Correctness and Verbosity. For Correctness, we classify responses as: (1) Correct, successfully addresses all satisfaction conditions with a working solution; (2) Partially Correct, addresses some conditions or requires minor modifications; or (3) Incorrect, fails to address the core issue or provides non-working solutions. For Verbosity, responses are classified as: (1) Appropriate, provides necessary information concisely; (2) Verbose, contains unnecessary details or repetitions; or (3) Terse, lacks sufficient explanation. These classifications are determined by Sonnet3.7 acting as an automated judge agent, which evaluates satisfaction condition fulfillment, execution results (when applicable), and compares with reference solutions (see Appendix A.5 for the prompt).
4.2.2 Sampling Strategy
Given the multi-turn nature of our evaluation framework, running all 3,286 issues through interactive simulation would be computationally prohibitive. Furthermore, issue counts vary significantly across languages and turn lengths, which could bias the evaluation outcomes. To address these concerns, we constructed a balanced, representative evaluation subset using a two-stage sampling strategy. For each language, we first selected up to five issues containing Dockerfiles (where available), and then sampled additional issues to ensure ten samples per turn-length: 1,2,3,4, and 5+. When specific ranges had fewer than ten available issues, the remaining quota was reallocated to the preceding bucket. Finally, we obtained 350 datasets for the all-time repositories (50 per language) and 194 datasets for the recent repositories (C: 18; C#: 25; C++: 28; Java: 8; JavaScript: 39; Python: 50; TypeScript: 26).
4.2.3 Quantitative Metrics - Correctness
Table 2 presents the performance metrics and conversation statistics for all evaluated models. We observe a consistent performance gap between recent and historical repositories. On recent repositories, correctness rates range from 7.22% (Haiku 3.5) to 16.49% (ChatGPT 4.1 Mini), while on all-time repositories, these rates increase substantially to 13.58-29.14%. ChatGPT 4.1 Mini demonstrates superior performance across both datasets, achieving the highest correctness (16.49% recent; 29.14% all-time) and lowest error rates on recent repositories (53.09%), while DeepSeek R1 shows the lowest error rate on all-time repositories (32.57%).
Figure 4: Side-by-side comparison of model responses to a Docker port-remapping issue: Haiku 3.5’s incomplete solution (middle) fails to address key requirements, while ChatGPT 4.1 Mini’s successful response (right) satisfies all three user conditions (highlighted in colored boxes).
4.2.4 Qualitative Example
Figure 4 illustrates a GitHub issue where the user asks how to remap ports for both the main application and the Cloudflare proxy to avoid conflicts with other apps. It presents two contrasting model stories: Haiku 3.5 provides an incomplete response, either omitting clear proxy-port guidance or introducing incorrect configurations, while ChatGPT 4.1 Mini offers a successful solution - clearly explaining how to change the main app’s port, noting that the proxy port doesn’t need to be exposed, and providing a conflict-free docker-compose.yml. A simulated user agent drives each LLM through a multi-turn dialogue, asking follow-up questions until all satisfaction conditions are met or a predefined turn limit is reached. Where applicable, the agent also validates the recommended configuration using recorded build outcomes. Once the conversation concludes, an automated LLM-Judge assigns a final label - Correct, Partially Correct, or Incorrect - based on comparisons with the original GitHub thread, execution results, and the predefined satisfaction conditions.
4.2.5 Conversation-Length Analysis
Table 2 compares the average turn counts between original GitHub threads and CAB-generated dialogues. When models produce correct answers, CAB conversations are similar in length to real GitHub threads - typically spanning 2–3 turns. In contrast, incorrect cases tend to result in longer CAB dialogues, often by 1–2 additional turns. This suggests that when the model-as-maintainer provides vague or incomplete responses, the simulated user - like a real developer - asks follow-up questions for clarification.
4.2.6 Language-Specific Analysis
Figure 6 reports the percentage of Correct responses - excluding partially correct answers - for each model across seven programming languages, split by recent and all-time GitHub repositories. The analysis reveals stark contrasts in language difficulty and model specialization.
Statically typed languages such as C#, C++, and Java remain particularly challenging, especially on recent issues. For instance, in the recent C# dataset, most models achieve less than 13% correctness. Similarly, correctness on recent C++ issues hovers below 15% for all models except ChatGPT 4.1 Mini. These results suggest models often struggle with precision and strict type constraints in newer repositories.
By contrast, dynamically typed languages like JavaScript and Python show relatively stronger performance on the all-time dataset. ChatGPT 4.1 Mini and Sonnet 3.7 Think both reach 44% correctness on JavaScript, while DeepSeek R1 and Sonnet 3.7 Think achieve over 30% on Python and TypeScript. Nevertheless, performance on recent repositories remains low across the board. Even for JavaScript, the best model - Sonnet 3.7 Think - achieves only 15.4% correctness. These results highlight the increased complexity or novelty of recent code issues.
Overall, Sonnet 3.7 Think (M6) consistently ranks among the top performers in JavaScript, TypeScript, and Python. ChatGPT 4.1 Mini (M1) also shows strength in all-time datasets for JavaScript and TypeScript. No model, however, maintains high correctness across both typed and dynamic languages, particularly in recent repositories.
4.3 Human Alignment Evaluation
Each GitHub issue was independently reviewed by three annotators. A total of 21 professional contractors participated in the evaluation, with an average of 5.1 years of programming experience. Annotators brought diverse technical backgrounds, spanning Python, JavaScript/TypeScript, C#, and Java, as well as familiarity with frameworks such as React, Django, Spring Boot, and .NET Core.
Overall, across 663 extracted conditions, 86.3% were judged accurate (i.e., correctly capturing what the user wanted), but only 65.7% were deemed complete (i.e., covering all of the user’s expectations). This indicates high precision - when a condition is extracted, it is usually correct - but lower recall, as annotators frequently identified missing satisfaction criteria. For example, in C++, 97.9% of extractions were accurate, yet 73.3% of cases required additional conditions. This suggests our pipeline favors conservative, reliable extractions over exhaustive coverage.
To put the 86.3% accuracy figure in context, we note that the average pairwise agreement among human raters was 82.4%, with some pairs achieving perfect agreement ($\kappa=1.0$) and others showing lower consistency. From the result, we interpret it as a strong result, indicating that the extracted conditions are nearly as reliable as those agreed upon by independent human annotators. Detailed annotation guidelines - including definitions of correctness and completeness, abstraction principles, and example outputs - are provided in Appendix G. Full per-language results and agreement statistics are available in Appendix F.
5 Limitations
While CodeAssistBench (CAB) offers a realistic and scalable framework for evaluating multi-turn programming assistance, it has several limitations. These include conservative condition extraction, limited evaluation coverage, templated user behavior, language scope constraints, and lack of statistical testing. We provide a full discussion in Appendix B.
6 Conclusion and Future Work
Our study introduces CodeAssistBench (CAB), a fully automated benchmark for evaluating multi-turn programming assistance grounded in real-world developer interactions. Unlike prior benchmarks that emphasize single-turn or synthetic tasks, CAB simulates realistic conversations over full codebases using containerized environments and issue-specific satisfaction criteria. Through extensive experiments, we find that state-of-the-art language models struggle with complex, multi-turn dialogues - particularly on recent repositories, where even the best models achieve only 16.49% correctness. CAB’s automated pipeline enables scalable dataset expansion while supporting rigorous, reproducible evaluation across languages, project types, and time periods, offering a more faithful measure of model capabilities than existing benchmarks.
Several promising directions could extend CAB’s impact, including improving satisfaction condition extraction to balance precision with recall, expanding language coverage beyond the current seven languages, incorporating more sophisticated user simulation strategies to better capture nuanced developer interactions, and developing specialized metrics for different assistance categories. We hope CAB facilitates deeper insights into the limitations of current systems and helps guide the development of future AI assistants that better support real-world software engineering workflows. We further discuss potential societal benefits and risks of CAB in Appendix C.
Appendix A LLM Filtering Prompts
A.1 Issue-Level Relevance Filter
To identify issues suitable for benchmarking, we use a 7-question prompt to assess resolution status, specificity, clarity, and safety. Below is the exact prompt provided to the LLM:
Please evaluate the following GitHub issue and its comments:
Title: {title}
Author: {author}
Body:
{body}
Comments:
{comments}
Based on this conversation, please answer the following questions with Yes or No:
1. Is the problem resolved by someone other than the author (not self-answered)?
2. Does the conversation contain confirmation from the author that the problem has been resolved?
3. Is the problem a specific technical issue (not a feature request, opinion, or open-ended question)?
4. Is there a clear, definitive solution provided within the conversation?
5. Can the solution be directly applied without requiring additional context or resources?
6. Does the conversation contain any personally identifiable information (PII) such as Email addresses, Phone numbers, Physical addresses, Full names (beyond just GitHub usernames), Passwords or credentials, Personal identification numbers, IP addresses, or Any other sensitive personal information?
7. Can this problem be reproduced and solved using the provided solution today (April 2025)?
Please provide your answers in the format:
1. [Yes/No]
2. [Yes/No]
3. [Yes/No]
4. [Yes/No]
5. [Yes/No]
6. [Yes/No]
7. [Yes/No]
The model responds with binary answers to each question, which we then parse to determine whether the issue is usable.
A.2 Message-Level Comment Filter
For message filtering, we construct the full conversation context and ask the LLM to identify non-contributory messages (e.g., "+1", "Thanks", "Bump"). A comment is retained unless explicitly marked for removal. The prompt follows:
Analyze each comment in this GitHub issue conversation:
{conversation}
Your task is to identify ONLY comments that have ABSOLUTELY NO support-related value.
A comment should ONLY be removed if it falls into ALL of these criteria:
- Contains NO technical information
- Provides NO context about the issue
- Asks NO relevant questions (technical or process-related)
- Provides NO status updates or next steps
- Offers NO feedback on proposed solutions
- Contains NO clarifications about the users situation or environment
- Has NO administrative or process value (like assigning work, requesting more info)
Examples of comments to remove:
1. Pure social messages at conversation end: "Thanks!", "Cool", "thumbs-up imoji"
2. Empty status updates: "+1", "Same issue", "Any updates?", "Bump" with no additional context
3. Completely off-topic discussions unrelated to the issue
IMPORTANT: Preserve comments that show the natural flow of support interaction. If a comment contains ANY support-related value, even if minimal or alongside thanks/acknowledgements, DO NOT remove it.
List ONLY the comment numbers that should be removed because they have absolutely no support-related value.
Format:
NUMBERS: <comma-separated list of numbers>
EXPLANATION: <specific reasons why these comments add no support-related value>
If no comments should be removed, respond with:
NUMBERS: none
EXPLANATION: All comments contain some support-related value or context
A.3 Automated Dockerfile Synthesis and Repair
To ensure that every benchmarked issue can be built and tested in a fully self-contained environment, we automatically (i) generate an initial set of Dockerfile candidates and (ii) iteratively refine any candidate that fails to build.
Phase 1: Candidate Generation.
For each GitHub issue we collect a rich context bundle—repository URL, commit SHA, truncated issue body, README, a structure summary of the repo, up to two GitHub Actions workflows, and (if available) a reference Dockerfile from the same project. We then prompt the LLM to produce 5 self-contained Dockerfiles, each of which must (1) install dependencies, (2) clone the repo at the specified commit, and (3) build the project without executing the user’s code. The system prompt fixes the LLM’s persona (“expert Docker engineer”), while the user prompt supplies the per-issue context. By requesting plain text only (no Markdown fences) we can pipe the response directly to docker build.
SYSTEM: You are an expert Docker engineer who creates Dockerfiles to build and validate GitHub issues.
Repository URL: {repo_url}
Title: {issue_title}
Commit SHA: {commit_sha}
Issue description (truncated to 3 kB):
{issue_body}
Context:
- README
- Repo structure summary
- GitHub workflow files (optional)
- Reference/Original Dockerfile (optional)
Create a Dockerfile that
1. installs all dependencies in the issue,
2. clones the repository and checks out {commit_sha} or user given project version,
3. builds the project (no test or run commands)
IMPORTANT: Return only the raw Dockerfile content - no Markdown, no commentary.
Listing 1: Prompt for candidate generation
Phase 2: Fault-Directed Repair.
If every candidate in Phase 1 fails to build, we capture the build log and feed it—together with the failing Dockerfile and the same context bundle—into a repair prompt. The LLM is asked to produce an improved Dockerfile that specifically addresses the observed errors. We repeat this loop for up to three attempts or until a build succeeds.
SYSTEM: You are an expert Docker engineer who specializes in fixing Dockerfiles that failed to build.
USER:
Repository URL: {repo_url}
Issue #: {issue_number} Title: {issue_title}
Commit SHA: {commit_sha}
Failing Dockerfile:
{candidate_dockerfile}
Build error (truncated to 3 kB):
{build_error}
Provide a corrected Dockerfile that
1. removes the above error(s),
2. keeps the minimal environment needed to build,
3. follows the same constraints as the generation prompt.
IMPORTANT: Return only the raw Dockerfile content - no Markdown, no commentary.
Listing 2: Prompt for candidate repair
A.4 Satisfaction-Condition Extraction
To evaluate whether an assistant’s reply actually meets a developer’s needs, we first distill each GitHub issue thread into a small set of user-satisfaction conditions—explicit criteria that any acceptable answer must fulfill. The extraction procedure is entirely automated and comprises two stages.
Stage 1: LLM-based Extraction.
We prompt a language model to extract these conditions from each conversation using two coordinated prompts: a system prompt that defines the task, abstraction level, and response format, and a user prompt that injects issue-specific content. The model returns a JSON object describing each condition along with a brief explanation.
You are an expert at analyzing GitHub issues and extracting user satisfaction conditions-the criteria by which any answer will be judged.
A satisfaction condition states WHAT the user needs, not HOW to implement it.
# Abstraction guide
TOO SPECIFIC - "Use numpy.where(...)"
GOOD LEVEL - "Vectorized conditional selection"
TOO GENERIC - "A working solution"
# Must-have properties
1. TRANSFERABLE (solution-agnostic)
2. VERIFIABLE (pass/fail is clear)
3. EVIDENCED (grounded in user utterances)
4. NEED-FOCUSED (problem, not implementation)
Return exactly this JSON:
"satisfaction_conditions": [
"condition": "...",
"explanation": "..."
Do not wrap the JSON in markdown fences.
Listing 3: System prompt for satisfaction-condition extraction
Given this GitHub conversation:
Title : {title}
Author : {author}
Question: {body}
Comments (chronological, at most 100):
{formatted_comments_json}
Extract every user satisfaction condition.
Remember: they are *criteria*, not solutions.
Output exactly one JSON object as described in the system prompt-no extra text.
Listing 4: User prompt for satisfaction-condition extraction
Stage 2: Post-hoc Verification.
The raw JSON returned by the LLM is parsed and each condition is passed through a lightweight verifier that checks:
  1. 1.
    the text parses as valid JSON;
  2. 2.
    each entry includes both a condition and an explanation;
  3. 3.
    the explanation quotes or paraphrases evidence found in the thread.
Conditions that fail verification are discarded. Issues with no surviving valid conditions are excluded from the benchmark.
A.5 LLM-Judge
To assess the quality of the maintainer’s answer in each conversation, we employ a structured LLM prompt that simulates a judge agent. The agent is provided with (i) the original GitHub issue (title, body, and comments), (ii) the set of user satisfaction conditions, (iii) the maintainer’s answer to be evaluated, and (iv) optional Docker validation results. The LLM is instructed to evaluate the answer along three axes: technical correctness, alignment with user satisfaction conditions, and verbosity.
Evaluation Prompt.
The judge agent uses the prompt shown in Listing 5:
You are a judge evaluating the maintainers answer to a users technical question.
Your task is to determine if the maintainers answer is:
1. TECHNICALLY CORRECT
2. SATISFIES USER CONDITIONS
3. APPROPRIATE VERBOSITY
IMPORTANT: For Docker-related issues:
- The answer must be technically correct AND
- The Docker build/test process must succeed
If Docker validation fails (Success: False), the answer is considered INCORRECT.
Provide your evaluation in this format:
TECHNICAL CORRECTNESS: [CORRECT / PARTIALLY CORRECT / INCORRECT]
ALIGNMENT SCORE: X/Y CONDITIONS MET (Z%)
CONDITION 1: [TRUE / FALSE] <brief description>
VERBOSITY ASSESSMENT: [TERSE / APPROPRIATE / VERBOSE]
VERDICT: [CORRECT / PARTIALLY CORRECT / INCORRECT]
KEY ISSUES:
- Issue 1
- Issue 2
REASONING:
Detailed explanation of correctness and alignment.
Listing 5: Prompt for LLM-Judge
Inputs.
The judge agent receives:
  • The full conversation (title, question body, comments),
  • User satisfaction conditions,
  • Maintainer’s generated answer,
  • Docker validation logs (if applicable).
Verdict Criteria.
The final judgment is based on:
  • Correct: Fully accurate and satisfies all user conditions.
  • Partially Correct: Minor technical flaws or partial condition satisfaction.
  • Incorrect: Major errors, unmet conditions, or failed Docker validation.
Post-Processing.
The LLM’s output is parsed to extract:
  • Technical correctness,
  • Number of conditions satisfied,
  • Verbosity assessment,
  • Final verdict,
  • Key issues and reasoning.
Appendix B Extended Discussion of Limitations
While CodeAssistBench (CAB) provides a realistic and scalable framework for evaluating multi-turn programming assistance, it has several limitations.
First, our satisfaction condition extraction prioritizes precision over recall. Although 86.3% of extracted conditions were judged accurate by human annotators, only 65.7% were complete, suggesting that models may be unfairly penalized for omitting criteria not fully captured by our automated pipeline.
Second, evaluation is performed on a sampled subset of 544 issues (from a pool of 3,286), constrained by computational cost. This may skew results away from rare or edge-case conversations.
Third, the simulated user uses BM25-matched historical responses to simulate follow-ups. While grounded in real-world interactions, this approach may underrepresent the full diversity of developer behaviors, especially in ambiguous or exploratory contexts.
Fourth, CAB only includes issues with successfully synthesized Docker environments. This excludes legacy or atypically configured projects, introducing a bias toward actively maintained, modern repositories with standard build systems.
Fifth, our evaluation targets only seven programming languages and repositories with permissive open-source licenses. The benchmark does not currently assess proprietary codebases, enterprise workflows, or other language ecosystems (e.g., Rust, Go, Swift).
Sixth, we fixed generation temperatures to promote reproducibility: temperature=0 for conversation responses and temperature=0.7 for Dockerfile generation (to encourage diversity across five sampled candidates). While this setup ensures consistency and comparability, it may not reflect each model’s optimal decoding parameters for task performance.
Finally, we do not report error bars or statistical significance tests. While aggregate model trends are clear, fine-grained comparisons—especially between close-performing models—should be interpreted cautiously.
We believe addressing these limitations—by improving condition recall, scaling human annotation, diversifying simulation strategies, and exploring a broader language and tooling spectrum—will further enhance CAB’s utility as a long-term benchmark for conversational coding agents.
Appendix C Broader Impacts
CodeAssistBench (CAB) has the potential to advance the development of safer and more capable AI programming assistants, particularly in realistic, multi-turn developer scenarios. By surfacing failure modes and limitations in current models, CAB encourages the creation of tools that better align with developer needs, which may enhance productivity and learning. However, there are also potential negative impacts. For instance, if developers rely too heavily on inaccurate AI suggestions in real-world settings, this could introduce subtle bugs or security issues. Additionally, the use of scraped GitHub data—even when filtered—may raise concerns around privacy or attribution if not carefully managed. We encourage responsible use of CAB, including proper consent for dataset extension and monitoring downstream usage of models trained or benchmarked using this framework.
Appendix D Hardware Requirements and Runtime Environment
Most stages of the CodeAssistBench (CAB) pipeline—including dataset generation, model evaluation, and human alignment analysis—are driven by API-based interactions with GitHub and commercial LLM endpoints. As a result, these stages require minimal local compute resources and can be executed efficiently on standard cloud instances or modest local machines.
The only hardware-intensive component is the Dockerfile synthesis and validation phase. This stage involves generating candidate Dockerfiles and building them in parallel across hundreds of repositories to ensure correctness and reproducibility. Due to the storage and memory demands of large-scale container builds, we recommend using machines equipped with at least 1TB of local storage and 32GB of memory.
All experiments were conducted on AWS g5.16xlarge instances, each with 1 NVIDIA A10G GPU, 64 vCPUs, and 256GB of RAM. While GPU acceleration was not required (as all model queries were served via commercial APIs), the instance’s large memory and CPU capacity were beneficial during Docker-based validation. The total estimated compute usage was approximately 100 CPU-hours, with dataset generation taking around 3 days and experimental runs completing in approximately 4 days.
Appendix E Supplementary Material
Figure 5: Distribution of GitHub issue conversation lengths by programming language. Each turn corresponds to a maintainer response; for example, a 1-turn conversation consists of a user question and a single maintainer reply, while longer conversations reflect additional back-and-forth exchanges.
Table 3: Docker environment build success rates by programming language
Language Environment-dependent Successful Builds Success Rate
C 17 17 100%
C# 30 21 70%
C++ 53 28 53%
Java 58 52 90%
JavaScript 34 33 97%
Python 90 77 86%
TypeScript 12 10 83%
Total 294 238 81%
Figure 6: Cumulative comparison of model performance across seven programming languages. Each bar represents a model’s prediction outcome distribution on GitHub issues, broken down into Correct (green), Partially Correct (orange), and Incorrect (red) responses. The top row corresponds to evaluations on recent repositories, while the bottom row shows results on all-time repositories. Models are labeled as M1–M6 in the following order: M1: ChatGPT 4.1 Mini, M2: DeepSeek R1, M3: Llama 3.3 70B, M4: Haiku 3.5, M5: Sonnet 3.7, M6: Sonnet 3.7 (Think Mode).
Figure 7: Cumulative verbosity distribution of model responses across seven programming languages. Each bar shows the proportion of responses classified as Appropriate (green), Verbose (orange), and Terse (gray, rarely observed). The top row corresponds to evaluations on recent repositories, while the bottom row shows results on all-time repositories. Models are labeled as M1–M6 in the following order: M1: ChatGPT 4.1 Mini, M2: DeepSeek R1, M3: Llama 3.3 70B, M4: Haiku 3.5, M5: Sonnet 3.7, M6: Sonnet 3.7 (Think Mode). Most models tend to generate verbose responses, with only a few models (notably M6) achieving higher rates of appropriately concise answers.
Appendix F Human Alignment Details
We conducted a targeted evaluation to assess the accuracy and completeness of automatically extracted user satisfaction conditions. We randomly sampled 10 GitHub issues per language across seven programming languages and collected annotations from three independent human raters. Each annotator evaluated two aspects: Condition Accuracy and Condition Completeness. Across 210 records, annotators reviewed a total of 663 satisfaction conditions. The detailed per-language breakdown is shown in Table 4.
Table 4: Human Alignment Evaluation: Accuracy, Recall, and Additional Condition Needs per Language
Language Total Conditions Correct Incorrect Accuracy Recall Additional Needed (%)
C 87 80 7 91.95% 96.67% 3.3%
C++ 96 94 2 97.92% 26.67% 73.3%
C# 96 68 28 70.83% 70.00% 30.0%
Java 102 87 15 85.29% 86.67% 13.3%
JavaScript 93 80 13 86.02% 53.33% 46.7%
Python 102 79 23 77.45% 66.67% 33.3%
TypeScript 87 84 3 96.55% 60.00% 40.0%
Overall 663 572 91 86.27% 65.71% 34.3%
Annotator Agreement.
We computed pairwise agreement and Cohen’s Kappa across the three annotators. The average percentage overlap was 82.39%, with some pairs reaching perfect agreement (100.00%, $\kappa=1.0$) and others showing low or negative $\kappa$, highlighting variability in interpretation and the need for further calibration in future rounds.
Appendix G Human Annotation Instructions
This section provides the full instructions shown to human annotators for evaluating the quality of user satisfaction conditions in CodeAssistBench (CAB). These instructions were designed to ensure consistency, transparency, and alignment with user intent during the evaluation process.
Ethics and Compensation
Annotators were external contractors who provided informed consent prior to participation. They were compensated at or above the local minimum wage, in accordance with the NeurIPS Code of Ethics. The task involved minimal risk, and no personally identifiable or sensitive information was collected.
Task Overview
This annotation task involves evaluating user satisfaction conditions automatically extracted from GitHub issue threads. A satisfaction condition describes **what** the user needs from the maintainer’s response—not **how** to implement it. These conditions should reflect the user’s goals, constraints, or expectations based on the full conversation.
Each annotator was asked to review 70 GitHub issues, with three annotators assigned per issue. A calibration phase with 10 examples was used to ensure alignment before the full task began.
Annotation Interface
Annotators were presented with the following information for each task:
  • Issue title and initial question
  • Full GitHub thread, including all comments and author roles
  • The list of model-generated satisfaction conditions
Annotator Instructions (Verbatim)
The following is the exact text shown to annotators:
Overview
This project focuses on verifying user satisfaction conditions automatically extracted from GitHub issue conversations using a language model. A satisfaction condition describes what the user needs from the maintainer’s response—not how it is implemented.
You will review GitHub issues and evaluate each proposed satisfaction condition for: - Correctness: Does the condition accurately reflect user needs? - Completeness: Are there missing conditions that should be included?
For each issue, you will be given: - The original GitHub title and user question - All follow-up comments (including roles and timestamps) - The model-generated satisfaction conditions
Your tasks: 1. For each listed condition: - Mark it as Correct or Incorrect - Provide a short justification (e.g., “User confirmed this need in a follow-up comment.”) 2. Suggest any missing conditions that should be added, with justification.
Annotation Format:
  "original_conditions": [
      "condition": "string",
      "judgment": "Correct" | "Incorrect",
      "justification": "string"
  "additional_conditions (if any)": [
      "condition": "string",
      "justification": "string"
Levels of Abstraction for Conditions:
Avoid overly specific or overly generic phrasing. Good conditions should reflect the what, not the how.
Too Specific (Avoid): - "Use numpy.where(condition, x, y)" - "Set max_depth=5 in the RandomForest constructor"
Good Abstraction Level: - "A vectorized approach to conditional element selection" - "Guidance on suitable tree depth settings"
Too Generic (Avoid): - "A working solution" - "Help with the problem"
Criteria for a Good Condition: - Transferable: Not tied to a specific implementation - Verifiable: Can be clearly judged as satisfied or not - Evidenced: Grounded in what the user said or implied - Needs-Focused: Describes the user’s goal/problem—not the method
Example Annotation
Below is an anonymized real example used during calibration:
Issue Title: “How can I play RTSP stream without audio codecs?”
User Post: How can I play RTSP stream without audio codecs? I need only video. I can’t start watching because the camera uses G.711 for audio.
Model-Generated Conditions:
  • Explanation of how the player handles unsupported audio codecs
  • Confirmation that video playback is possible without audio codec support
Annotator Output:
"original_conditions": [
"condition": "Explanation of how the player handles unsupported audio codecs",
"judgment": "Correct",
"justification": "User asked about G.711 and the maintainer explained it would be dropped automatically."
"condition": "Confirmation that video playback is possible without audio codec support",
"judgment": "Correct",
"justification": "The maintainer confirmed that video-only playback is supported."
"additional_conditions": []
Quality Control
Each annotator completed a 10-example calibration round. Only those with consistent agreement advanced to the full task. Inter-annotator agreement statistics are reported in Appendix F.
Annotator Backgrounds
A total of 21 annotators participated in the human evaluation task. All annotators had at least 2 years of programming experience (mean: 5.1 years), with backgrounds spanning industry roles in software engineering, quality assurance, and backend development.
  • Languages covered: JavaScript/TypeScript, Python, Java, C#
  • Common frameworks: React, Angular, Spring Boot, Django, .NET Core
  • Cloud & infra skills: AWS, SQL, HTML/CSS
  • Review process: Each GitHub issue was annotated by 3 independent reviewers.
A full table of individual annotator experience is available upon request but omitted from the appendix to preserve readability.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The Abstract and Section 1 clearly articulate CAB’s three core contributions: (1) an automated dataset generation pipeline, (2) a realistic multi-turn Q&A dataset grounded in GitHub issues, and (3) a chat-driven evaluation framework. These claims are consistently supported by the methods and experiments presented throughout the paper, ensuring alignment between stated contributions and actual work.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations in Section 5, including issues such as conservative satisfaction condition extraction (favoring precision over recall), limited evaluation coverage due to sampling, and simulation artifacts from templated user behavior. These limitations help contextualize the scope and applicability of our benchmark and findings.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This work is entirely empirical and does not present formal theorems or proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We detail the dataset generation pipeline, evaluation setup, and sampling strategy in Section 3 and Section 4. All code, data, prompts, and instructions needed to reproduce the main experimental results are publicly available at the URL provided in the Abstract.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: As stated in the Abstract, the complete CAB benchmark, evaluation framework, and dataset are available at https://anonymous.4open.science/r/CAB-CBA3/. The repository includes instructions for accessing the raw and processed data, build environments, and running the evaluation pipeline.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe the evaluation subset construction—including sampling by language, turn length, and Docker presence—in Section 4. Model generation parameters, including fixed temperatures (0 for dialogue and 0.7 for Dockerfile generation), are specified in Appendix B. The full evaluation setup, prompts, and execution logs are available at https://anonymous.4open.science/r/CAB-CBA3/.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Due to the multi-turn, agentic nature of CAB’s evaluation—where each model response triggers follow-up interactions and environment-level validation—each data point is expensive to generate. Running multiple trials per issue (e.g., with varied sampling or decoding) was not feasible at scale. Nevertheless, the evaluation spans over 500 diverse issues across 7 languages and multiple turn lengths, offering reliable performance trends despite the lack of formal confidence intervals.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The paper provides compute details in Appendix D, including hardware specifications, resource requirements, and an estimate of total compute usage.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We use only publicly available GitHub issues and apply LLM-based filtering for personally identifiable information (PII), as described in Appendix A.1, to ensure that no sensitive personal data is included in our dataset or analysis.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We include a brief note in the Conclusion directing readers to Appendix C, where we discuss CAB’s potential societal benefits—such as improving developer productivity and transparency in LLM-based tools—and risks, including overreliance on automated suggestions and possible misuse in security-sensitive environments.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This work does not release any high-risk assets such as generative models or sensitive datasets; thus, safeguards are not applicable.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: In Section 3.1.1, we restrict repository selection to those under permissive licenses, including MIT, Apache-2.0, ISC, Zlib, CC-BY, and CC0-1.0, to ensure license compliance. Original repository authors are credited in our metadata, which includes license names and source URLs.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The CAB dataset’s structure and components are described in Section 3. Documentation including usage instructions, license, and known limitations is provided at https://anonymous.4open.science/r/CAB-CBA3/.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: The paper includes a complete description of the annotation task and protocol in Appendix G, including the full text of instructions given to annotators, example inputs and outputs, abstraction guidelines, and JSON formatting schema. Annotators were external contractors who provided informed consent and were compensated at or above the local minimum wage, in accordance with the NeurIPS Code of Ethics.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: The human annotation study was reviewed and approved through our organization’s internal human subjects review process. Annotators were external contractors who provided informed consent prior to participation.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We explicitly detail the use of Sonnet 3.7 for message filtering, environment generation, and satisfaction condition extraction in Sections 3.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.